Find the value of the x. Y is 30 degrees and side is 14


1
Expert's answer
2021-01-27T03:42:48-0500


1) x =14 * "cos(30\degree)" = 14 * "\sqrt{3}" = 7 * "\sqrt{3}"

2) x = "sin(30\degree)"/14 = (1/2)/14 = 1/28

3) x = 14/"sin(30\degree)" = 14/(1/2) = 28

4) x = 14/"cos(30\degree)" = 14/("\sqrt{3}"/2) = 28/"\sqrt{3}"

5) x = 14/"cot(30\degree)" = 14/"\sqrt{3}"

6) x = 14 * "cot(30\degree)" = 14 * "\sqrt{3}"